3. A golfer, putting on a green, require three strokes to " hole the ball." During the first putt, the ball rolls 5.0 m due east. For the second putt, the ball travels 2.1 m at an angle of 20.0 degree north of east. The third putt is 0.50 m due north. What displacement(magnitude and direction relative to due east) would have been needed to "hole the ball" on the very first putt?
Answer #1

East components

5 + 2.1cos20 + 0 = 6.973 m

North components

0 + 2.1sin20 +0.50 = 1.218 m

d = √(6.973 ² + 1.218²)

d = 7.08 m

tanθ = 1.218 / 6.973

θ = 9.91° N of E